OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x - 2y = 8.

OpenStudy (anonymous):

The Choices are x intercept = (8, 0); y intercept = (0, 4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (-8, 0); y intercept = (0, -4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (aravindg):

put x=0 solve for y .You get y intercept put y=0 solve for x you get x intercept

OpenStudy (aravindg):

Hope that helps

OpenStudy (anonymous):

Not really... I'm still so confused

OpenStudy (aravindg):

x - 2y = 8. x=0 -2y=8 y=?

OpenStudy (aravindg):

x - 2y = 8. y=0 x=8

OpenStudy (aravindg):

can you do it now ?

OpenStudy (anonymous):

So we have the equation; x - 2y = 8 So, first let's find the x-int; To find the x-int, we need to set y = 0 and solve for x. x - 2(0) = 8 x - 0 = 8 x = 8 So the -int is (8, 0) Now to find the y-int we need to replace x with 0 and solve for y, using what I did above, how would you start that?

OpenStudy (anonymous):

It would be 0,-4

OpenStudy (anonymous):

Yes.

OpenStudy (aravindg):

now u also got x=8 for x intercept !

OpenStudy (aravindg):

So the final answer will be option b